# PSAppendix - Selected Primary Studies

### v1

@@ Abstract.
With the increasing trend of Machine Learning (ML) enabled software applications, the paradigm of ML Operations (MLOps) has gained tremendous attention of researchers and practitioners. MLOps encompasses the practices and technologies for streamlining the resources and monitoring needs of operationalizing ML models. Software development practitioners need access to the detailed and easily understandable knowledge of MLOps workflows, practices, challenges and solutions to effectively and efficiently support the adoption of MLOps. Whilst the academic and industry literature on the MLOps has been growing rapidly, there have been relatively a few attempts at systematically synthesizing and analyzing the vast amount of existing literature of MLOps for improving ease of access and understanding. We conducted a Multivocal Literature Review (MLR) of 150 relevant academic studies and 48 gray literature to provide a comprehensive body of knowledge on MLOps. Through this MLR, we identified the emerging MLOps practices, adoption challenges and solutions related to various areas, including development and operation of complex pipelines, managing production at scale, managing artifacts, and ensuring quality, security, governance, and ethical aspects. We also report the socio-technical aspect of MLOps relating to diverse roles involved and collaboration practices across them through the MLOps lifecycle. We assert that this MLR provides valuable insights to researchers and practitioners seeking to navigate the rapidly evolving landscape of MLOps. We also identify the open issues that need to be addressed in order to advance the current state-of-the-art of MLOps.
@@
$c$s: Software and its engineering Software development methodsccs: General and reference Empirical studiesccs: General and reference Surveys and overviews
@@ 1. Introduction
Currently, there is a huge trend of incorporating Artificial Intelligence (AI) and Machine Learning (ML) components into software systems to take advantage of their well established inferential and predictive capabilities. These systems, known as ML-enabled systems, provide enhanced business value through data-driven decisions. The increasing demand for AI/ML-enabled solutions and the availability of novel technologies/tools to support the engineering of such systems have increased AI/ML adoption. IBM Global AI Adoption Index for 2022 reports that AI/ML adoption is increasing rapidly with 44% of the organizations working on embedding AI/ML in their products or processes in 2022 (ibm, 2022). However, the productionalization of the AI/ML components, which is the process of transforming experimental ML models into an ML-enabled software product operating in a production environment remains one of the main challenges. The majority of the organizations that build ML pilots fail to deploy their models in production or fail to maintain successful operation of deployed models (Salama et al., 2021) due to novel operationalizing challenges of ML-enabled software systems compared to traditional software. These challenges include ML-specific technical debt introduced during model development, the need for continuous monitoring, model retraining and delivery due to the data-dependant probabilistic nature of ML systems, collaboration among different actors such as data scientists, software engineers, and software operators, the critical need of ensuring responsible and trustworthy AI, and the complexity of integration and scaling of ML-enabled systems (Salama et al., 2021; ibm, 2022). To tackle these challenges, MLOps has emerged as a paradigm that introduces practices, processes, techniques and technologies for enabling rapid and reliable productionalization of ML systems.
MLOps was first proposed by the research community in 2015 as a solution to speed up the ML lifecycle management and to achieve the high level of scalability expected by business applications (Data and Team, 2023; Sculley et al., 2015) by streamlining time and resources required to operationalize ML models. Similar to DevOps, the concept of MLOps has emerged from production challenges, driven by the industry, and encompasses a wide range of practices, techniques, and technologies. For instance, the prominent aspect of MLOps is the definition and management of an automated pipeline, such as for testing and deploying ML-enabled software to production. Such pipeline can be considered the codification of the repetitive yet hidden processes between teams responsible for the life cycle of ML-enabled software. Automating these processes could potentially reduce human errors and increase their velocity. Moreover, the act of defining pipelines could be catalytic to facilitate the communication between teams, and to help identify bottlenecks and pitfalls. Another practice often cited as related to MLOps is the establishment and utilization of shared version-control repositories for datasets, features, and models to provide a consistent source of truth for the automation pipelines and increase the visibility across teams to reduce duplicate work. Another relevant practice under MLOps umbrella is on-demand monitoring services, which could initiate and increase the fidelity of feedback loops. When connected to the pipeline, the improved monitoring could lead to an increase in velocity. Research shows that proper operationalization of ML-enabled systems depends on utilizing such robust MLOps practices, processes, and technologies (Salama et al., 2021). According to the survey conducted by Databricks, while the percentage of ML models making it to production is still low, it sees a considerable increase with the rapid advancements in MLOps solutions (e.g., MLFlow, Huggingface, SageMaker, Google Vertex AI). As of January 2023, 34% of the experimental models are registered as candidates for production which is a significant increase compared to 20% in 2022 (PS, 114).
The apparent success of MLOps in bridging the gap between development and deployment of ML-enabled software systems has sparked a lot of interest among both researchers and practitioners, who have been increasingly reporting different types of MLOps practices and tools for supporting MLOps. As building ML systems is yet not as mature as traditional software systems and due to their inherent complexity, different organizations tend to follow a variety of approaches for developing and operationalizing ML systems based on their unique use cases and organizational structure. MLOps spans across a broad spectrum of activities (starting from data collection to model maintenance), a plethora of practices, tools and technologies supporting and automating each activity and interdisciplinary collaborations (data scientists, ML engineers, software engineers, MLOps engineers). Such inherent complexity and diversity make it harder for practitioners and researchers to navigate current MLOps landscape, which hinders effective adoption and further improvements of MLOps practices and solutions. Despite the increasing amount of the peer-reviewed and gray literature on MLOps, there is significant confusion surrounding MLOps due to it’s complexity, diversity, and broad scope (PS, 154).
Hence, it is important to systematically select, analyze and organize the MLOps literature covering it’s various aspects such as how MLOps is being perceived, MLOps practices and technologies, and MLOps adoption challenges and the related solutions to overcome the reported adoption challenges. Whilst there have been a few attempts to review the State-Of-The-Art (SOTA) of MLOps such as reported in these papers (Kreuzberger et al., 2023; Lima et al., 2022; Kolltveit and Li, 2022; Faubel and Schmid, 2023; Mboweni et al., 2022), these studies have reviewed only academic literature and do not cover several important aspects of MLOps. Given MLOps is quite applied paradigm about which there is an increasing amount of gray literature, mainly published by industry. Moreover, the peer-reviewed literature on MLOps has also grown significantly since the previous reviews (Kreuzberger et al., 2023; Lima et al., 2022; Kolltveit and Li, 2022; Faubel and Schmid, 2023; Mboweni et al., 2022), there is an important and urgent need of a comprehensive review of the peer-reviewed and gray literature for building an evidence-based body of knowledge on MLOPs.
We conducted a Systematic Multivocal Literature Review (MLR) covering peer-reviewed and gray literature on MLOps published up to September 2023. This study aimed at systematically selecting, synthesizing and analyzing the state-of-the-art MLOps perception, best practices, adoption challenges and solutions for building a body of knowledge to be used by researchers and practitioners interested in MLOps.
The main contributions of our study, linked to our Research Questions (RQs), are as follows:
  • It conducts a MLR for building an evidence-based body of knowledge of MLOps from peer-reviewed and gray literature.
  • It helps improve our understanding of how MLOps is perceived by both researchers and practitioners, consisting of MLOps definitions (RQ1.1), activities (RQ1.2), roles, and responsibilities (RQ1.3).
  • It systematically identifies and analyzes the state-of-the-art best practices (i.e., processes, techniques, and technologies) reported to support MLOps (RQ2).
  • It determines the challenges faced by practitioners during MLOps adoption and derives the future research directions for improving MLOps adoption (RQ3).
2. Related Literature Reviews
We consider existing systematic review studies related to MLOps and conduct a qualitative comparison with our work in terms of the type of the considered primary studies, research questions covered in the review study, number of primary studies used in the review, and considered time period for primary study selection. We identified a total number of 8 related studies, consisting of three Multivocal Literature Reviews (MLR) (i.e., these studies collected data from both academic and gray literature), and five Systematic Literature Reviews (SLR) (i.e., they solely collected data from academic literature). Table 1 presents the comparison of these related reviews with our work to highlight our contributions.
Table 1. Comparison of Related Work
Study Is multivocal RQ1.1 RQ1.2 RQ1.3 RQ2 RQ3 No. Primary Studies Timeframe
[…]
(Lima et al., 2022) No $Circle \Circle \Circle$ 30 A Up to Jul, 2021
[…]
Our work Yes $CIRCLE \CIRCLE \CIRCLE \CIRCLE$ 150 A, 60 G 2013 - Sep, 2023
[…]
A: Academic literature study, G: gray literature study
Kreuzberger et al. (Kreuzberger et al., 2023) reported a mixed-method research, where they conducted an interview with practitioners in addition to a SLR. Their work targeted one RQ: “what is MLOps”, which they answered with an overview of MLOps activities, roles and responsibilities and used these insights to create a unifying definition for MLOps. While they also discussed several open challenges related to adopting MLOps, it is not the main objective of the study. Lima et al. (Lima et al., 2022) reported a SLR, where their RQs aimed to identify activities related to ML model deployment, MLOps related roles and responsibilities, operationalization tools for ML models, and challenges in ML model deployment. However, this study lacks comprehensive data synthesis to provide indepth analysis of roles and responsibilities. Additionally, their discussion on MLOps related technologies is limited to MLOps platforms extracted from academic literature and their discussion on challenges does not elaborate on the challenges related to MLOps adoption. Kolltveit et al. (Kolltveit and Li, 2022) reported an SLR on the operationalization of ML systems, which covers all steps following model training and evaluation. Specifically, they answered four RQs focused on tooling and infrastructure related aspects of MLOps such as; how are ML models operationalized, what are the main challenges in operationalizing ML models, what tools and software infrastructure are used to operationalize ML models, what are the feature gaps in the tooling and infrastructure used to operationalize ML models. However, since the focus is only on tooling and infrastructure for operationalizing ML models, this work does not provide a comprehensive view of MLOps best practices considering the entire ML lifecycle. Faubel et al. (Faubel and Schmid, 2023) reported an SLR on the technical overview of MLOps. Specifically, the study presented MLOps definitions and listed the most common activities, and listed tools for automating activities in MLOps. However, this review does not consider roles and task distributions in MLOps, thus lacking emphasis on socio-technical aspects in MLOps. Mboweni et al. (Mboweni et al., 2022) reported an SLR on the state-of-the-art DevOps practices in ML-enabled systems. Specifically, this work aimed to report how MLOps is understood and differs from DevOps. The reported content provides a higher-level overview of MLOps definitions, presents key roles of MLOps and associated tasks, and describes MLOps activities by constructing an MLOps architecture covering ML lifecycle.
Steidl et al. (Steidl et al., 2023) reported a mixed-method investigation (i.e., SLR and a semi-structured review study) focusing specifically on continuous development pipelines of AI models, which is one aspect of MLOps practices. This work focused on analyzing definitions of various terms related to continuous development pipelines for AI such as DevOps for AI, CI/CD for AI, MLOPs, etc., tasks handled by these pipelines, and the challenges related to implementation, adaption and usage of pipelines for continuous development of AI. John et al. (John et al., 2021) delivered a MLR on the adoption of MLOps in practice. More specifically, the authors derived a framework for MLOps that shows the activities that take part in the continuous development of ML models and maps the practices of companies to MLOps maturity model to show how practices evolve through stages. However, they provided a high-level picture rather than elaborating on details regarding these questions. Lwakatare et al. (Lwakatare et al., 2020) presented an MLR on the RQ of “how to enable modern software development process with continuous delivery for AI-enabled systems”. The authors considered one academic and eight gray literature items. The study discusses tools and techniques used in the continuous delivery of ML systems. Important details are available concerning the development stages of ML-enabled systems.
Our analysis of the secondary studies presented in Table 1 demonstrates, that no existing work contributes to building a holistic and timely depiction of MLOps landscape by answering all three research questions. Moreover, the majority of the related work relies solely on the academic literature (Kreuzberger et al., 2023; Lima et al., 2022; Kolltveit and Li, 2022; Faubel and Schmid, 2023; Mboweni et al., 2022). Given that MLOps is a rapidly evolving, practice-oriented field, insights from gray literature sources such as industry white papers and company websites offer valuable and timely perspectives on how MLOps is perceived by the community and adopted in real-life settings. Furthermore, the rapid improvements in MLOps call for a systematic review of the most recent literature to provide a timely view of the MLOps landscape. However, coverage by existing reviews extends only until 2022.
Considering these, our MLR aims to provide a thorough understanding of MLOps encompassing its definition, tasks, roles, state-of-the-art best practices and adoption challenges. We combine theoretical analysis from 150 academic studies with more practical knowledge from 48 gray literature items published up to September 2023. Also, we characterize MLOps as covering emerging practices such as socio-technical practices, complex pipeline management, and ethical and governance considerations. Additionally, we discuss future directions in detail which facilitates further advancement in the field.
@@ 3. Methodology
We conducted a Multivocal Literature Review (MLR) to provide a comprehensive overview of MLOps in terms of its definition, activities, involved people and their responsibilities, state-of-the-art best practices for implementing MLOps, and MLOps adoption challenges faced by the practitioners, mapping them to solutions proposed in the literature.
An MLR study allows researchers to collect and present information from the industry when the formal/academic literature is not sufficient (Garousi et al., 2019). Since MLOps is an emerging topic in the field of AI and software engineering, only considering academic publications is not sufficient to cover all the knowledge related to this topic. Industry practitioners have been implementing and getting experience in MLOps to integrate ML-enabled systems into their software development pipelines, their knowledge is a crucial element in understanding MLOps. Thus, gray literature such as industrial white papers, blog posts, and company websites, are the key sources for expanding the knowledge obtained from academic literature. Therefore, we believe an MLR would be the most suitable methodology for the MLOps topic. We performed our MLR following the guidelines reported by Garousi et al. (Garousi et al., 2019).
Consequently, our MLR consists of three main stages: 1) planning the review, 2) conducting the review, and 3) reporting the review. The planning stage includes identifying the need for the review, specifying the RQs, developing a review protocol containing search strings for the peer-reviewed and gray literature, and defining inclusion/exclusion criteria for primary study selection. In the second stage, we conducted the search for peer-reviewed and gray literature and selected the primary studies. Afterward, data extraction and synthesis were performed following Thematic Analysis guidelines proposed by Braun et al. (Braun and Clarke, 2006). In the final stage, we have organized and reported the synthesized data. Figure 1 represents our MLR methodology.
@@
3.1. Research Questions
Our work addresses the RQs listed in Table 2 to provide a comprehensive overview of the current MLOps landscape from academic and gray literature.
@@ 3.1. Research Questions
RQ1 - How is MLOps perceived by researchers and practitioners? RQ1 focuses on collecting, synthesizing, and reporting an overview of MLOps and its core principles that would reflect the views of both researchers and practitioners under 3 sub-RQs.
RQ1.1 - How is MLOps defined in the literature? RQ1.1 aims to provide insights on common and varying definitions of MLOps to facilitate better understanding of its foundational aspects to contribute towards creating a standardized usage by the community.
RQ1.2 - What are the primary activities and tasks of MLOps? RQ1.2 reports the state-of-the-art activities that take part in MLOps to help practitioners adopt MLOps in their contexts more efficiently and allow the readers to gain a clear business value/ perspective.
RQ1.3 - What are the practitioner roles and responsibilities in MLOps? RQ1.3 highlights the responsibilities of people who take part in MLOps settings. Establishing a common understanding of roles and their responsibilities leads to more productive collaboration practices within the complex MLOps team structure.
RQ2 - What are the state-of-the-art practices and techniques proposed for MLOps? A complete view of best practices, processes and supporting technologies is required for companies to plan for successful implementation of MLOps by minimizing the technical debt and improving the maintainability of the ML systems as they grow in complexity over time. To this end, RQ2 aims to provide state-of-the-art MLOps practices and techniques covering social and technical aspects across the entire life cycle of ML models.
RQ3 - What are the challenges practitioners face when adopting MLOps? The motivation behind this question is to gain a deeper understanding of the challenges practitioners face when implementing MLOps to see the bottlenecks in MLOps adoption and identify solutions proposed in the current literature. Addressing the state-of-the-art challenges would be beneficial to practitioners to overcome those when they adopt MLOps for productionalizing their ML-enabled software systems. Further, existing challenges often lead to new and highly impactful research directions and innovative solutions.
Table 2. Research Questions addressed in this MLR
3.2. Study Search and Selection Process
We based our search process on the guidelines provided by Garousi et al. (Garousi et al., 2019) for conducting MLR. We aimed at retrieving the maximum number of academic, peer-reviewed, and gray literature related to the research questions. Our search strategy consists of three main steps: automatic search method using inclusive search strings, primary study selection by applying quality assessment and snowballing to identify further studies (Wohlin, 2014).

### v2

@@ Abstract.
MLOps has emerged as a key solution to address many socio-technical challenges of bringing ML models to production, such as integrating ML models with non-ML software, continuous monitoring, maintenance, and retraining of deployed models. Despite the utility of MLOps, an integrated body of knowledge regarding MLOps remains elusive because of its extensive scope due to the diversity of ML productionalization challenges it addresses. Whilst the existing literature reviews provide valuable snapshots of specific practices, tools, and research prototypes related to MLOps at various times, they focus on particular facets of MLOps, thus fail to offer a comprehensive and invariant framework that can weave these perspectives into a unified understanding of MLOps. This paper presents a Multivocal Literature Review that systematically analyzes a corpus of 150 peer-reviewed and 48 grey literature to synthesize a unified conceptualization of MLOps and develop a snapshot of its best practices, adoption challenges, and solutions.
@@
journal: CSURjournalvolume: 1journalnumber: 1article: 1publicationmonth: 1ccs: Software and its engineering Software development methodsccs: General and reference Empirical studiesccs: General and reference Surveys and overviews
@@ 1. Introduction
There is a huge trend of incorporating artificial intelligence (AI) and machine learning (ML) components into software systems to construct ML-enabled systems and take advantage of their well-established inferential and predictive capabilities. IBM Global AI Adoption Index for 2022 reports that AI/ML adoption is increasing rapidly, with 44% of the organizations working on embedding AI/ML in their products or processes in 2022 (ibm, 2022). As of January 2023, 34% of the experimental models are registered as candidates for production, which is a significant increase compared to 20% in 2022 (PS, 114).
Despite the interest in ML-enabled systems, most organizations that build ML pilots fail to bring their ML models to production (Salama et al., 2021) due to novel operationalizing challenges of ML-enabled software systems compared to traditional software. Productionalization of ML models refers to the process of transitioning ML models from a laboratory setting to a production environment, where they must handle live traffic from real users and provide inference services reliably at scale while ensuring safety. Successful productionalization of ML models requires organizations to address multiple challenges, both technical and social in nature. For instance, one prominent issue is the integration of ML models into the surrounding software systems [143]. To enable this integration, software engineers must be able to package the trained model [143], [138], [141], [125], [139], [156] and implement the necessary data transformation logic to feed data from the surrounding software components to the ML model with correct syntax and semantics [179]. Failure to do so could result in erroneous predictions, such as when feeding metric measurements to a model that expects imperial measurements. The details regarding syntax, semantics, and data transformation used by ML models might not be apparent or readily documented. Consequently, resolving this technical challenge requires software engineers to collaborate closely with data scientists and ML engineers despite these disciplines’ differing languages and cultures [151].
Another common problem in the productionalization of ML models lies in the monitoring and maintaining ML-enabled software in production [107]. The IT operations team must implement additional telemetry to monitor the behaviour of ML models, allowing them to detect when models suffer from concept drifts or violate safety constraints [152]. Again, this task requires a skill set different from that typically found among IT operations teams. It necessitates collaboration between IT staff and data scientists with distinct disciplinary cultures and lexicon [151].
Additionally, retraining and updating models when problems such as drifts are detected in production represents another challenge in operating ML-enabled systems. The retraining process necessitates new data, likely sourced from the production environment. This data may contain sensitive personal information and must undergo appropriate preprocessing [152]. Developing this data pipeline requires collaboration between software engineers, IT operations, data engineers, and data scientists. During the model training phase, it’s critical to maintain provenance records of experiments for analysis and compliance purposes [116], [165]. After a new model is created, software engineers and IT operators face the challenge of rolling out the new models quickly and safely and need to rely on strategies like A/B testing, canary deployments, and blue/green deployment [111], [168]. Velocity and safety ultimately drive the success of ML model productionalization.
The concept of MLOps emerged as a key solution to address these challenges of ML model productionalization. MLOps was first proposed by the research community in 2015 to expedite ML lifecycle management while achieving the high scalability required by business applications (PS, 115; Sculley et al., 2015). Research indicates that the effective operationalization of ML-enabled systems is contingent on employing robust MLOps practices, processes, and technologies (Salama et al., 2021). A survey conducted by Databricks in 2023 indicates that, although the percentage of ML models reaching production remains low, there has been a significant increase with the rapid advancements in MLOps solutions (e.g., MLFlow, Hugging Face, SageMaker, Google Vertex AI).
Despite the widespread prevalence of MLOps, an evidence-based body of knowledge regarding MLOps practices and tools grounded in a unified conceptualization of the MLOps concept remains elusive for several reasons. First, the sheer scope of MLOps encompasses a broad array of practices, techniques, and corresponding technologies, complicating the effort to systematize the knowledge. For example, MLOps can be viewed from the perspective of ML “operations” (i.e., bringing ML models from the lab into deployed software systems), leading to research related to the packaging of ML models, interoperability with surrounding software systems, and testing of ML-enabled systems (e.g., [179], [172], [156], [128], [168]). MLOps is also often conceptualised as ”DevOps for ML”, leading to the development of practices and tools for defining and managing automated pipelines of ML lifecycle activities (e.g., [136], [172]). Additionally, MLOps is also viewed from the perspective of systematic management of ML experiments and related assets (e.g., datasets, feature sets, trained models), leading to practices such as applying version control to ML assets and recording experiments to support both analysis and provenance for compliance and security (e.g.,(PS, 130, 164)). On-demand and continuous monitoring of model performance and safety in a production environment is yet another critical practice associated with MLOps (e.g.,(PS, 131)). Finally, given that ML model productionalization involves collaboration among multiple teams, many MLOps best practices emphasize the need for organizational and cultural changes to facilitate necessary cross-team collaboration in handling ML productionalization (e.g., establishing project teams [116], [111], building specialized ML engineering teams).
The diversity of information sources on MLOps compounds the issue. Particularly, gray literature such as training materials and best practices from various platform providers [116], [165], [154], [106], [107] play a significant role in MLOps since it is inherently driven by practice. This gray literature must be compared and contrasted with peer-reviewed literature to synthesise a comprehensive and unbiased understanding of the field, which requires substantial time and effort. This challenge is exacerbated by the previously mentioned diversity in what constitutes “MLOps”. There is considerable confusion surrounding MLOps due to its complexity, diversity, and broad scope (PS, 154), creating obstacles for both practitioners and researchers navigating the current MLOps landscape, hindering effective adoption and subsequent improvements in MLOps practices and solutions.
In this paper, our objective is to build an evidence-based knowledge base for MLOps based on a unified conceptualization of the MLOps concept by integrating input from the gray literature and peer-reviewed literature to address the challenges that practitioners and researchers face when working with MLOps. The unified conceptualization of MLOps in this context denotes a coherent and comprehensive model that accounts for the various points of view on MLOps with respect to its definition, activities carried out within its scope, and the roles and responsibilities involved. Such a conceptualization creates a framework for identifying and classifying the diverse practices, processes, and technologies under the MLOps umbrella. “Evidence-based” in this context signifies that the knowledge base and conceptualization will be synthesized inductively from the literature. This knowledge base’s benefit is providing practitioners with a framework for understanding, assessing, and adopting MLOps practices while offering researchers a roadmap for open research opportunities. We applied the Systematic Multivocal Literature Review (MLR) and thematic analysis methodologies to address the research goal. The MLR methodology enables us to systematically identify, select, and extract information from peer-reviewed and gray literature on MLOps published up to September 2023. The thematic analysis allows us to systematically analyze the extracted data and synthesize the conceptualization of MLOps, best practices, adoption challenges, and solutions to build a robust body of knowledge for researchers and practitioners interested in MLOps.
The multifaceted nature of MLOps and the abundance of primary studies classified under this umbrella have led to many literature reviews exploring various aspects of MLOps. For example, Steidl et al. (Steidl et al., 2024) conducted a tertiary analysis that identified 33 reviews encompassing 1,397 primary studies on MLOps and found minimal overlap in the coverage of these reviews. Whilst the existing literature reviews provide valuable snapshots of specific practices, tools, and research prototypes related to MLOps at various times, they focus on particular facets of MLOps (e.g., the DevOps angle, the experiment management angle, the continuous monitoring angle), thus fail to offer a comprehensive and invariant framework that can weave these perspectives into a unified understanding of MLOps. Such a framework is crucial because it aligns current and future practices and tools with their challenges. Although snapshots of MLOps technologies and practices can quickly become outdated as tools evolve, a robust framework allows for analyzing and categorizing present and future developments, paving the way for creating new snapshots. Thus, the systematic synthesis of a conceptual framework for MLOps from both peer-reviewed and gray literature represents our study’s core focus and contribution (Section 3). Building upon this foundation, we identify and categorize the state-of-the-art practices and techniques for implementing MLOps, forming a ”snapshot” of the literature (Section 4). Finally, we highlight practitioners’ challenges when adopting MLOps and identify future research directions (Section 5).
@@ 2. Methodology
A Multivocal Literature Review study (MLR) allows researchers to collect and present information from the industry when the formal/academic literature is not sufficient (Garousi et al., 2019). Since MLOps is an emerging and practitioner-driven topic, only considering academic publications is not sufficient to cover all the knowledge related to this topic. Industry practitioners have been implementing and getting experience in MLOps to integrate ML-enabled systems into their software development pipelines, their knowledge is a crucial element in understanding MLOps. Thus, gray literature such as industrial white papers, blog posts, and company websites, are the key sources for expanding the knowledge obtained from academic literature. Therefore, we believe an MLR would be the most suitable methodology.
We performed our MLR following the guidelines reported by Garousi et al. (Garousi et al., 2019). Consequently, our MLR consists of three main stages: (1) planning the review, (2) conducting the review, and (3) reporting the review. In the planning stage, we defined the research questions and developed the protocol for identifying and selecting literature. In the conducting phase, we followed the specified protocol to identify and choose literature, extract data from the literature, and synthesize the extracted data to answer the research questions. We employed the Thematic Analysis methodology based on the guidelines proposed by Braun et al. (Braun and Clarke, 2006) to carry out data synthesis. In the final stage, we write up the synthesis results. Figure 1 represents our MLR methodology.
@@
2.1. Planning the Review
We designed three main research questions (RQs) to systematically explore MLOps through academic and gray literature to address our research goals: (i) establish a foundational understanding of how MLOps is defined and perceived, (ii) identify and map current practices, and (iii) identify challenges and solutions within the field. Table 1 outlines the RQs in detail and the motivations behind each.
Table 1. Research Questions addressed in this MLR
@@ 2.1. Planning the Review
RQ1 - Conceptualization of MLOps Exploring foundational aspects of MLOps is important for creating a common and standardized understanding of MLOps, optimizing workflows, enabling efficient adoption of MLOps practices across diverse contexts, and fostering collaboration.
RQ1.1 - How is MLOps defined in the literature? RQ1.1 aims to provide insights on common and varying definitions of MLOps to facilitate better understanding of its foundational aspects to contribute towards creating a standardized and consistent usage within the community.
RQ1.2 - What are the primary activities and tasks that define an MLOps workflow? RQ1.2 identifies the core components of MLOps, focusing on the essential activities and tasks within an MLOps workflow. Outlining these key elements, provides a foundation for process optimization and standardization, while offering practitioners actionable insights for efficient adoption and clarifying the operational impact of MLOps in their specific contexts.
RQ1.3 - What are the practitioner roles and responsibilities in MLOps? RQ1.3 highlights the responsibilities of people who take part in MLOps settings. Establishing a common understanding of roles and their responsibilities leads to more efficient workflows and productive collaboration practices within the complex MLOps team structure.
RQ2 - What are the state-of-the-art best practices and techniques for implementing MLOps? A complete view of how MLOps activities and tasks are performed in practice covering, frameworks, processes, tools and best practices to implement them is required for companies to plan for successful implementation of MLOps by minimizing the technical debt and improving the maintainability of the ML systems as they scale and grow in complexity over time. To this end, RQ2 aims to provide state-of-the-art MLOps practices and techniques covering social and technical aspects across the entire life cycle of ML models.
RQ3 - What are the challenges practitioners face when adopting MLOps, and what solutions have been proposed to address them? Gaining a deeper understanding of the challenges practitioners face when implementing MLOps along with proposed solutions in the current literature, is crucial to see the bottlenecks in MLOps adoption. RQ3 addresses the state-of-the-art challenges that would be beneficial to practitioners to overcome those when they adopt MLOps for productionalizing their ML-enabled software systems. Further, existing challenges often lead to new and highly impactful research directions and innovative solutions.
2.2. Study Search and Selection Protocol
We based our protocol on the guidelines provided by Garousi et al. (Garousi et al., 2019) for conducting MLR. Our protocol consists of three main steps: automatic search method using inclusive search strings, primary study selection by applying quality assessment and snowballing to identify further studies (Wohlin, 2014).